the sum of the ages of a mother and daughter is 75 and the difference in their ages is 29 years. how old is each?
Answer by ikleyn(52790)

.
the sum of the ages of a mother and daughter is 75 and the difference in their ages is 29 years. how old is each?
-----------------------------------------------------------

Let m = mother's age, d = daughter age. Then

m + d = 75,
m - d = 29.

Add these equation. You will get 

2m = 75 + 29 = 104.

What to do next?

Solve yourselve the problem from this point.